OpenStudy (anonymous):

If f(x) = x - 4, what does f(3) represent? The value of (x - 4) when x = 3 The value if x when (x - 4) = 3 The value of 3f(x) when x = 3 The value of f(x) when (x - 4) = 3

OpenStudy (texaschic101):

f(x) = x - 4 f(3) = 3 - 4 The value of (x - 4) when x = 3.....I think

OpenStudy (anonymous):

I have a similar question to this could you help me on that too?

OpenStudy (texaschic101):

I am not the best at functions...but I can look at it and see

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (texaschic101):

y = mx + b slope(m) = -3 (2,-5)..x = 2 and y = -5 now we sub...and since we know m, the slope, we need to find b, the y intercept -5 = -3(2) + b -5 = -6 + b -5 + 6 = b 1 = b so our equation is : y = -3x + 1

OpenStudy (texaschic101):

I would stay longer and help more, but there is somewhere I have to be in a little bit. I will be back on later.

OpenStudy (anonymous):

Thank you!!!

OpenStudy (texaschic101):

anytime :)